OpenStudy (marceljr59):

solve for x: 2/3 (x − 2) = 4x.

OpenStudy (raffle_snaffle):

okay so we have \[\frac{ 2 }{ 3 }\left( x-2 \right)=\]

OpenStudy (raffle_snaffle):

4x lol

OpenStudy (raffle_snaffle):

 distribute the 2/3 through.